OpenStudy (anonymous):

A car is traveling a constant velocity. After two hours the car has traveled 220 miles. Assuming a linear equation fits this situation, find the distance that the car will travel after after five hours.

OpenStudy (anonymous):

\(v=d/t\) find the speed first using d=220 miles, t=2 hrs. then using the same formula, calculate for the d, using the speed that you found and t=5 hrs

OpenStudy (anonymous):

Okay, v=220/2 v= 110 110 x 5? or 110/5?

OpenStudy (anonymous):

\(\LARGE {v=\frac{d}{t}\\d=\frac{d}{v}}\)

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

velocity is constant and he is right v= d/t

OpenStudy (anonymous):

so you have d = 220 and t = 2 now you have to find d at t = 5

OpenStudy (anonymous):

so 220/2 = d/5

OpenStudy (anonymous):

so d = 220*5/2 :) calculate

OpenStudy (anonymous):

550

OpenStudy (anonymous):

yep